### v1

# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation

Abstract.
The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in the task of mobile UI automation. However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined ones, are unscalable and unfaithful. To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device agent execution and faithful, scalable agent evaluation. By observing that the task execution process only transfers UI states, LlamaTouch employs a novel evaluation approach that only assesses whether an agent traverses all manually annotated, essential application/system states. LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with real mobile environments for task completion. (2) Fine-grained UI component annotation that merges pixel-level screenshots and textual screen hierarchies to explicitly identify and precisely annotate essential UI components with a rich set of designed annotation primitives. (3) A multi-level state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen with unpredictable UI layout/content dynamics. LlamaTouch currently incorporates four mobile agents and 495 UI automation tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones for more diverse mobile applications. Evaluation results demonstrate the LlamaTouch’s high faithfulness of evaluation in real environments and its better scalability than human validation. LlamaTouch also enables easy task annotation and integration of new mobile agents. Code and dataset are publicly available at https://github.com/LlamaTouch/LlamaTouch.
copyright: none
1. Introduction
Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious, cumbersome smartphone operations. These agents are particularly beneficial for individuals with visual or hand impairments, or in situations where using a screen is not practical (e.g., driving). Notable mobile agents, such as Apple Siri (Apple, 2024) and Google Assistant (Google, 2024b), have become indispensable services on smartphones. The recent advent of large language models (LLMs) and multimodal LLMs have facilitated researchers building more powerful mobile agents (Li et al., 2024; Wen et al., 2023; Wang et al., 2023; Lee et al., 2023). The key capability of those agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI automation, e.g., “forward the last email from Bob to Alice”.
Despite claims of powerful UI automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed. Unlike traditional machine learning models evaluated on well-established, static datasets, mobile agents need to interact with the dynamic, indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs. Simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
In general, there are two methods to evaluate mobile UI automation tasks, but neither achieves both high faithfulness and scalability. (1) The most intuitive approach is to request humans to verify the completion of tasks. However, human evaluation is difficult to reproduce (Chiang and Lee, 2023). The requisite human effort increases with the number of agents, tasks, and evaluation platforms. (2) The most popular approach used in most prior arts (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations. The key idea is to ask annotators to generate a correct sequence of actions that succeed on the task as data labels, and then compare agent-generated actions to these labels. Although this approach allows for some error tolerance, e.g., the variations in click positions on the screen (Rawles et al., 2023), it cannot cover all possible and “infinite” paths to complete a UI automation task. Consequently, it leads to a significantly higher false negative rate. For example, for task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged. Only taking one of these execution paths as the reference may incorrectly verify a task that is essentially completed. Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI automation capabilities, yet is impossible to evaluate in a static dataset. These limitations are further demonstrated in $\S$§$2.$
This paper presents LlamaTouch, the first testbed for evaluating mobile agents on real-world mobile environments without compromising faithfulness and scalability. The key idea of LlamaTouch is to check the task execution trace against a few “essential states” identified by the annotators, rather than matching them against predefined action sequences in static traces. For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened; and (2) the application is on the login page. Other operations, like app installation, are considered non-essential and should be ignored. During task execution, LlamaTouch enables mobile agents to retrieve only the task description from the static dataset, while device states are directly acquired from real-world devices. Actions produced by mobile agents are executed on those smartphones, and all UI interaction data are recorded as task execution traces. In the evaluation phase, LlamaTouch compares task execution traces with annotated essential states to determine whether a task has been completed.
To ensure faithful and scalable evaluation, LlamaTouch integrates two effective methods. (1) LlamaTouch adopts a fine-grained labeling mechanism for essential state annotation at both the screen level and single UI component level. It combines pixel-level screenshots and textual screen hierarchies to explicitly highlight important UI components. With a rich set of annotation primitives provided by LlamaTouch, it reduces human efforts to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the content of a textbox should be exactly matched. These annotated primitives are subsequently used for faithful evaluation. (2) During the evaluation process, LlamaTouch employs a multi-level state match algorithm that combines fuzzy and exact matches on diverse annotated UI states. It uses (i) approximate screen match that enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state match that detects and matches critical on-screen information.
Dataset and testbed. We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world Android environments. This dataset includes 495 distinct tasks encompassing a wide array of popular Android applications. We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with real Android environments. This testbed provides a collection of concise, widely used APIs, ensuring compatibility with most mobile agents. Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI automation tasks in real-world scenarios.
Evaluation. We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) as actual smartphones. Currently, LlamaTouch has four built-in agents, AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2023), and CoCo-Agent (Ma et al., 2024), along with 495 diverse tasks. The number of supported agents and tasks will be extended in the future. With human validation as the ground truth for task completion, LlamaTouch achieves nearly 80% accuracy in detecting completed tasks in real-world environments, while exact action match on static datasets fails to do so. We also reveal the limitations of current mobile agents in handling tasks practically in real-world environments.
Contributions are summarized as follows.
  • We observed the weakness of high false negative rates in evaluating mobile UI automation agents using static datasets. To address this, we proposed an evaluation design that only compares essential states rather than concrete action sequences.
  • We devised a method for annotating essential states using a variety of annotation primitives. This approach combines visually intuitive screenshots with semantically precise VHs to enable fine-grained and accurate UI component localization and annotation.
  • We designed a novel task evaluation system that employs both exact and fuzzy matching at various UI state levels. It enables faithful evaluation of mobile agents and scales well to dynamic execution environments.
  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI automation agents in real-world mobile environments. It comprises 495 UI automation tasks with human-annotated essential states. Four agents integrated in LlamaTouch demonstrate its faithfulness and scalability in UI automation task evaluation.
2. Background and Motivation
2.1. Agents for Mobile UI Automation
Mobile agents have simplified the cumbersome and dull operations on smartphones for users. The progression of mobile agents for mobile UI task automation can be categorized into three phases. (1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through pre-defined application programming interfaces. This approach is reliable while limited in structured and predictable tasks. (2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities were still confined by their training data. (3) Recently, LLMs and multi-modal LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2023; Wang et al., 2023). These models, owing to their vast knowledge base, can understand complex, real-world scenarios. Mobile agents powered by these models are able to interpret natural language instructions accurately and translate them into actionable tasks on smartphones’ screens. This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
Mobile agents typically operate with the following components. Controller is the brain of mobile agents. It interprets task instructions and UI contexts, and generates actions to be executed on the current UI context. The widely-used controllers include DL models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (GPT-3.5, GPT-4) (Wen et al., 2023; Wang et al., 2023; Lee et al., 2023), and multi-modal LLMs (GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
Input: UI representation. Existing mobile agents take a task description and a UI representation as the input of their controller. There are two basic types of UI representations: screenshots and VH. A screenshot is a visual capture of the current screen. A text-format VH provides a tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents. On top of screenshots and VH, some controllers further extract UI semantics to enhance UI understanding. For example, Yan et al. (Yan et al., 2023) overlies numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) also converts screenshots to bounding boxes and labels, enabling them comprehensible by LLMs. Further processing based on VH, like converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2023).
Output: actions. The output of controllers are actions to be executed on the current screen, such as click, swipe, and input text. Action parameters can be abstracted at different levels depending on the agent’s design and their input format. (1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023). It operates as a direct interaction with the screen like human operations. (2) Icon marker (Yan et al., 2023; Yang et al., 2023). The output target will specify a specific icon or graphical element within the UI representation. (3) HTML index (Wang et al., 2023; Wen et al., 2023; Lee et al., 2023). By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
Table 1. The comparison between mobile agent benchmarks. LlamaTouch is the first testbed designed for mobile agents driven by essential state match. LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of diverse UI states.
Benchmark Platform
Real-world
Tasks
Real-env
Task Exec
Fine-grained
UI Annotation
Essential
State Match
Rico (Deka et al., 2017) Mobile
PixelHelp (Li et al., 2020)
AndroidEnv (Toyama et al., 2021)
META-GUI (Sun et al., 2022)
MoTIF (Burns et al., 2022)
AITW (Rawles et al., 2023)
Mobile-Env (Zhang et al., 2023)
AndroidArena (Xing et al., 2024)
WebArena (Zhou et al., 2023) Web
LlamaTouch Mobile
Figure 1. Two major limitations of evaluating mobile agents on static datasets.
$2.$ Mobile UI Automation Benchmarks
As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI automation tasks, while none of them achieve both faithfulness and scalability.
Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets (task descriptions and UI representations). Mobile agents predict concrete actions on static UI representations, which will be compared with the ground-truth actions in datasets. While this approach is straightforward, it is insufficient to reveal the performance of mobile agents for two reasons. (1) Inaccurate exact action match are pervasive when a functionally correct action is deemed incorrect due to different action parameters. Figure 1(a) shows two cases. First, for click actions, the clickable area defined in the dataset may be narrow. While in reality, the area might encompass the entire UI component (marked with the red bounding box). Second, non-identical text inputs can lead to the same correct result (marked with green bounding boxes) in most search tasks. This issue has also been observed in previous literature but has not been solved (Yan et al., 2023; Rawles et al., 2023). (2) Lack of tolerance for different execution paths. In real-world environments, an UI automation task can usually be completed in various paths, as demonstrated in Figure 1(b1) and Figure 1(b2). However, pre-defined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
There are also work that enable agent execution in real-world environments. AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023) enable the execution of mobile automation tasks in real environments. However, they do not essentially support essential state match during the end-to-end task execution, therefore compromising evaluation accuracy. AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets does not fully tolerant redundant actions in task execution paths. They proposed a subsequence-based action match: a task is treated as completed when it contains the ground-truth action sequence as its subsequence. We take AndroidArena as a baseline in $\S$§4.3 to compare its evaluation accuracy with LlamaTouch. WebArena (Zhou et al., 2023) provides a real playground for Web agents. It uses essential states to evaluate task completion (e.g., the final result should be or should include some key information). LlamaTouch differs from WebArena on the mobile platform in both the essential state annotation and evaluation process: (i) LlamaTouch combines the visual screenshot with textual view hierarchy of the same screen for fine-grained and precise UI component identification; (ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
Human validation is usually used to validate whether a UI automation task is completed (Rawles et al., 2023; Taeb et al., 2023). However, the cost of human validation is too high, thereby scaling poorly to multiple tasks, agents, and mobile devices. LlamaTouch ensures high scalability as evaluating on static datasets while preserving faithfulness as human validation.
Figure 2. The workflow of on-device task execution.
3. LlamaTouch Design
Compared to previous evaluation approaches, LlamaTouch exhibits the following benefits.
$\bullet$ Practical on-device task execution ($\S$§3.1). Most of the previous evaluation methods are simulated on mobile UI interaction datasets. LlamaTouch, on the other side, enables mobile agent evaluation on real mobile devices, revealing their real capabilities in the wild.
$\bullet$ Fine-grained essential application state annotation ($\S$§3.2). By observing UI automation task execution only transfers essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched during task execution.
$\bullet$ Faithful and scalable task evaluation ($\S$§3.3). LlamaTouch evaluates mobile agents by comparing task execution traces with annotated application states. By combining exact matching with fuzzy matching algorithms on different application states, LlamaTouch achieves faithful and scalable task evaluation.
3.1. On-device Task Execution
Evaluating mobile agents on pre-defined traces, as discussed in Section 2.2, depicts significant inaccuracy. LlamaTouch empowers on-device UI automation task execution to reveal the real capabilities of mobile agents. To simplify this process, LlamaTouch operates as a bridge between existing mobile agents and real mobile devices (e.g., smartphones, Android emulators, and cloud device farms). As shown in Figure 2, mobile agents execute tasks on LlamaTouch by taking the following processes. 1 Mobile agents request one task instruction from LlamaTouch. The instruction comes from one supported dataset by LlamaTouch, e.g., AITW. 2 With this task instruction, mobile agents will then request UI representation (e.g., screenshot, VH) from LlamaTouch, which will be 3 forwarded to mobile devices. 4 Taking the task instruction and UI representation as inputs, mobile agents will predict an action to be performed on the current UI, and dispatch predicted actions to LlamaTouch. 5 LlamaTouch forwards and executes the agent-predicted action to real mobile devices. The processes from 2 to 5 are repeatedly executed until mobile agents treat the task as completed. During task execution, all UI representations and corresponding actions are captured for further evaluation in $\S$§$3.$ These UI representations (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application at the foreground of each screen; and (4) actions performed on every screen. Essential system states such as the list of installed applications will also be dumped for faithful mobile agent evaluation, as we will illustrate in $\S$§3.2. Appendix A.1 details APIs provided by LlamaTouch for mobile agent integration.
Table 2. Primitives of essential states in LlamaTouch that are used in the annotation and evaluation processes.
Match Type State Type Primitive Keyword Use Case
Fuzzy match UI state Screen info fuzzy<-1> Check if the contents on two screens are approximately identical.
Textbox fuzzy<n>
Check if the content of the target textbox is semantically similar
to the content of the original textbox<n> in the ground-truth UI.
Exact match Activity activity
A coarse-grained approach to determining if two UIs represent
the same functional screen in an application.
UI component
exact<n>,
exclude<n>
Check if the UI component is exactly identical to the UI component<n>,
or does not occur, in the ground-truth UI.
System state (Un)installation
installed<app>,
uninstalled<app>
Check if the target application named "app" has been successfully
installed/uninstalled.
Action Action
click<n>,
type<input_text>
Check if two actions and their parameters are identical.
Figure 3. Annotated essential state for task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes that should be exactly matched. The essential state represents the application state after task execution: “the shopping cart on bestbuy is empty”.
3.2. Essential Application State Annotation
Two major cases demonstrated in $\S$§2.2 make exact matching on pre-defined traces unfaithful. First, it demands that the agent-generated actions and their parameters match exactly with those in the dataset. Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, therefore unable to evaluate alternative task execution paths.
Insight: the task execution process transfers identifiable application states. The internal of the execution process of a UI automation task is changing application states that some of them can be explicitly represented by UI components. Even if the task execution paths differ, there are overlaps in the essential application states. We can utilize these overlapping application states to determine whether a task achieves some milestones or is completed. As the example shown in Figure 1(b), by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page. Other actions, like detecting whether the application is installed, can be omitted during evaluation.
To do so, essential states should be accurately identified and annotated. However, simply annotating application states at the whole UI representation level (e.g., a screenshot) and comparing screen-level similarity (Feiz et al., 2022), are far coarse-grained and may lead to inaccurate matching. For example, the screen contents of a web-shopping application could be subtly different due to slight movement or dynamically loaded contents across different executions. For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation (i.e., screenshot) to separate UI components. LlamaTouch simplifies the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components. These extracted UI components are combined with visually intuitive screenshots to provide precise overlayed bounding boxes and unique identifiers to annotators. Figure 3 shows an example of the VH-enhanced screenshot provided to annotators.
Primitives. LlamaTouch incorporates a list of primitives for essential state annotation. These primitives are used to extract essential information on the screen. Annotators are instructed to clearly identify and use primitives to represent application states that are informative and deterministic to validate the task execution results. Table 2 comprehensively shows these primitives and their use cases. Overall, LlamaTouch incorporates six types of primitives and nine keywords for annotation. These primitives can be divided into three types according to the state type they represent.
$\bullet$ UI state. The purpose of UI state annotation is to extract and compare whether two screens contain identical or similar information. As a screen may contain different types of components, LlamaTouch has four major primitives to cover all these UI components. (1) Screen info and (2) textbox are used to compare whether the whole screen or a dedicated textbox has similar contents. Screen info is primarily used for checking whether two UI representations are within the same screen (i.e., the same page of an application). Textbox primitive is used to annotate textboxes that may contain dynamic contents, such as a search box in a web-shopping application or the URL of a website. (3) The functionality of activity is like screen info: it can be used to approximately detect whether two UI representations are on the same screen of an application. (4) UI component is used to deterministically annotate the state of a UI component, including textbox, button, image, etc. For example, if the content of a textbox should be identical, it should be annotated with the exact keyword. States that require certain UI components should not occur on the screen should be annotated with the exclude keyword.
$\bullet$ System state. LlamaTouch involves system state annotation mainly inspired by the tasks in a previously built dataset: AITW (Rawles et al., 2023). For example, a task “install app YouTube Kids” can be detected directly using shell commands (e.g., pm list packages), without involving redundant UI operations. Currently, LlamaTouch supports programmatically checking application installation. The keywords can be easily extended when other system states are required to be detected.
$\bullet$ Action. Although the principle of annotation is to detect application state transfer during task execution. Sometimes, action on a specific UI representation is necessary to validate agent behaviour, especially when there is not enough state identifier shown in UI. LlamaTouch provides click and type keywords that covers most common actions.
Case study. Essential state annotation can be divided into the following processes. First, for each UI representation, LlamaTouch overlayed all functional UI components with numeric markers to the screenshot as shown in Figure 3. This is done by extracting the precise metadata of each UI component from textual VH. Then, annotators will identify what are essential states that should be checked during evaluation to ensure the task is completed. LlamaTouch simplify this process by only requiring annotators to explicitly identify what UI element should be matched. For example, after emptying the shopping cart on bestbuy, the screen will display a textbox with the content “Your cart is empty”. This textbox should be treated as an essential state, which is highlighted in a bounding box with numeric id 13 in the last screen of Figure 3, as it represents the state after task completion. As we anticipate the content of the textbox should be exactly matched, the annotated keyword is exact<13>. The keyword exact<27> is used to validate whether two screens are both in the shopping cart of bestbuy. All annotations along with the UI representations and task descriptions construct the essential state-powered dataset. These essential states are used by LlamaTouch for faithful task evaluation.
$3.$ Faithful and Scalable Task Evaluation
Annotated essential states ($\S$§3.2) and task execution traces captured on real devices ($\S$§3.1) are used for evaluation on LlamaTouch. LlamaTouch iterates essential states to determine whether a task execution trace goes through all annotated states sequentially. If so, this task is deemed completed. To achieve faithful evaluation, the most significant challenge lies in how to ensure the task execution trace matches the essential states. Task execution traces captured from real-world environments may contain dynamic screen content; it is vital to adapt the annotated essential states on a static dataset to dynamic, real-world contents, while achieving precise match on only critical information. To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components, to ensure faithful evaluation. It first approximately matches two screens according to their UI representation and activity. Then, within two matched screens, it compares the annotated primitives that should be fuzzily or exactly matched.
Approximate screen match is used to ensure two screens are on the same functionality page of an application, even under high screen content dynamics. This is the prior process for in-screen exact/fuzzy content match. To achieve this, LlamaTouch utilizes two annotation primitives proposed in $\S$§3.2: activity and screen info.
1. Application activity match. Activity represents an entry point for users to interact with the application (and, 2024). This is an obvious identifier to indicate whether two screens are in the same application. Screens with distinct functionalities typically have unique activity names, even if they are within the same application. For example, the activity of the main setting page is com.android.settings.Settings, while the value is com.wifiadmin.settings.WifiSettingsActivity in the Wi-Fi setting page. Typically, exact application activity match acts as a foundational filtering process to identify whether two screens are on the same page. However, some specific application design philosophies may make different functional pages of an application contain the same activity name (sin, 2018). Under this circumstance, LlamaTouch involve screen info for accurate screen match.
2. Fuzzy screen info match is utilized when the activity name cannot reveal different functional pages of an application. To better compare two screens, the primary thing is to only extract critical information from the screen. In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation as in prior work (Wen et al., 2023; Wang et al., 2023), while containing the types of every single UI component. LlamaTouch compares two simplified HTML representations of two screens with a cosine similarity of their sentence embeddings. If the result falls within a threshold, LlamaTouch treats two screens are matched. Two screens are similar when they are above a predefined threshold, e.g., 0.85 in our experiments. The fuzzy screen info match design helps LlamaTouch maintain faithfulness during evaluation when dealing with dynamic screen contents.
Mixed UI state match will be applied to matched screens after the approximate screen match process. A mix of annotated UI states will be checked in this phase, including both fuzzy match and exact match on UI components, actions, and system states.
1. Exact UI component match requires the annotated UI component should be identical in two screens to be compared, including all their attributes such as class, text, and selected in VH. This is especially useful for evaluating the content of a textbox, the status of a button (e.g., checked or not), or a selectable icon on the screen. Given a target UI component to be exactly matched, LlamaTouch will iterate nodes in VH of the matched screen, until one matched UI component is found. Exact UI component match fails if there is no matched node on available screens. In our dataset, exact UI component match contributes to 50.5% (688 out of 1,362) of annotated essential states.
2. Fuzzy textbox match is crucial for comparing content inside a textbox on the screen, especially when the content may be slightly different. Semantically similar searching keywords with the same intent that complies with a specific task description should be matched, as they will lead to the same results. For example, Figure 1(b) shows searching “Microsoft Excel” and “Excel” in the Google Play Store both display the target application. LlamaTouch extracts the content of the annotated textbox, and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
3. Action match. Although the initiative of LlamaTouch is to detect state transfer during task execution. There are still cases for evaluating concrete actions performed on the screen, such as clicking one specific UI component, or typing the correct captcha. LlamaTouch directly compares actions and their parameters of the annotated ones with their counterparts on the matched screens.
4. System state match is usually more efficient and accurate than merely comparing UI states for specific tasks that involve deterministic system states, e.g., installed applications. LlamaTouch currently supports checking whether an application is installed or not. Such system state is recorded during task execution in real-world mobile devices ($\S$§3.1). LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace.
Through the multi-level state matching algorithm, LlamaTouch achieves high evaluation accuracy on real-world task execution traces, while preserving the scalability of evaluating on static datasets. Experiments in $\S$§4 show the faithful evaluation of LlamaTouch.
4. Evaluation
4.1. Dataset
Table 3. LlamaTouch’s dataset statistics.
Task Category # Tasks # Apps Avg. Task Steps
Tasks from AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
LlamaTouch Generated 393 46 5.67 (3-42)
Total 495 57 7.01 (2-42)
LlamaTouch uses a combination of tasks in the previous AITW dataset (Rawles et al., 2023) and self-constructed ones for evaluation. Table 3 shows the statistics of this dataset. We sampled 102 tasks from AITW (Rawles et al., 2023), one of the most popular datasets for mobile device control. By observing the limited number of evaluated applications, we generated 395 tasks in 46 applications. Overall, this dataset includes 495 tasks, covering 57 popular Android applications with different difficulty levels (measured by the number of steps to finish a task). Every single data entry consists of three types of data.
  • A task description in natural language, e.g., empty the shopping cart on bestbuy.
  • A sequence of raw UI representations when recording these tasks. Each raw UI representation contains a screenshot, VH, activity name, and action performed on the current screen.
  • The annotated essential states by human annotators.
4.2. Experiment Setup
Mobile environment. LlamaTouch primarily utilizes an x86-64 Android emulator (Google, 2023) as the mobile environment for task execution. The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server with Ubuntu 18.04 OS. Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) will be tested on a real Google Pixel 5 device with the Android 14 OS.
Mobile agents. We selected four mobile agents that cover diverse types of brains, including supervised learning models, LLM, and large multi-modality models.
  • AutoUI (Zhan and Zhang, 2023) that employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and t5 (Raffel et al., 2020) with its variants as the brain for decision making. It takes pixel-level screenshots and task descriptions as the input.
  • AutoDroid (Wen et al., 2023) that uses LLMs for device control. It first takes textual screen VH and simplifies it to explicit, readable HTML representation. Then, the simplified HTML representation will be processed by LLMs and generate the corresponding action. We selected GPT-4 (gpt-4-0125-preview) (Achiam et al., 2023) as the backend of AutoDroid.
  • AppAgent (Yang et al., 2023) that utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch. In our evaluation, we used the version AppAgent without document pre-exploration for simplicity.
  • CoCo-Agent (Ma et al., 2024) that uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
All actions generated by these agents are applied to LlamaTouch’s Android environment through one adaption layer.
Evaluation methodologies. We used the following methodologies for evaluating whether an agent completes a task.
  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023). It compares the agent-generated action sequences in real-world environments with ground-truth action sequences in the dataset. When two action sequences are identical, the task is treated as completed. We utilize the action match algorithm in AITW (Rawles et al., 2023): two actions are matched only if they have identical action types and parameters.
  • Longest common subsequence (LCS)-based action match is proposed in AndroidArena (Xing et al., 2024). It is built upon step-wise action match by tolerating redundant actions between ground-truth ones. Two action sequences are matched when the execution action sequence contains the ground-truth sequence as the subsequence.
  • LlamaTouch that compares agent execution traces with our annotated essential states to check whether a task is completed.
  • Human validates whether a task is completed based on the agent execution traces. Humans are instructed not only to focus on the action (and its parameter) on each UI representation, but also the whole application state transfer during task execution. The results of human validation are treated as ground truth for comparing the accuracy of the three evaluation approaches.
Metrics. We primarily compare the (1) end-to-end task completion rate (TCR) of different agents and (2) the accuracy of different evaluation approaches, taking human validation as the ground-truth result. In step-wise action match and LCS-based action match, a task is completed only when the two action sequences are identical or the ground-truth one serves as the subsequence of the agent-generated action sequence, respectively. In LlamaTouch, a task is treated as completed when the agent execution trace passes through all essential states. For accuracy, a task is accurately evaluated only if the evaluation result matches the human validation result. The accuracy of an evaluation approach denotes the proportion of tasks correctly evaluated out of all tasks.
4.3. Task Completion Rate and Accuracy
Table 4. End-to-end task completion rate (TCR %) and accuracy (Acc. %) of different evaluation approaches of all tasks.
Mobile Agent
Step-wise
action match
LCS action
match
LlamaTouch Human
TCR Acc. TCR Acc. TCR Acc. TCR
AutoUI 0.00 98.18 0.00 98.18 4.44 96.57 1.82
AutoDroid 0.00 85.98 0.00 85.98 14.84 91.87 14.02
AppAgent 0.00 93.33 0.61 93.13 10.91 94.95 6.67
CoCo-Agent 0.00 97.97 0.00 97.97 4.47 96.34 2.03
Average 0.00 93.86 0.15 93.81 8.67 94.93 6.14
Table 4 presents the end-to-end TCR and accuracy of different evaluation designs. All three approaches achieve more than 90% accuracy on average. However, the step-wise action match and LCS-based action match fall short in recognizing tasks correctly executed by mobile agents, resulting in nearly 0% TCR. The high accuracy rates of these two approaches are attributed to the large portion of incomplete tasks; only 6% of tasks are deemed completed upon human validation. Compared to action match on static datasets, the TCR evaluated by LlamaTouch is 8.67%, which closely aligns with the results of human validation.
To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results that pervasively occurred in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks. Table 5 displays the number of tasks completed by agents and the evaluation accuracy for these tasks. Overall, agents successfully completed 30 tasks on average. Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy. This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets. Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods. The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
Table 5. Accuracy (Acc. %) of different evaluation approaches among all successful tasks in human validation.
Mobile Agent
Step-wise
action match
LCS action
match
LlamaTouch Human
Acc. Acc. Acc. # success tasks
AutoUI 0.00 0.00 77.78 9
AutoDroid 0.00 0.00 73.91 69
AppAgent 0.00 3.03 93.94 33
CoCo-Agent 0.00 0.00 70.00 10
Average 0.00 0.76 78.91 30
$4.$ Ablation Study
Table 6. Ablation study on different essential state primitives.
Evaluation design All tasks AITW tasks LlamaTouch tasks
TCR Acc. TCR Acc. TCR Acc.
Complete LlamaTouch 8.67 94.93 17.46 89.43 6.38 96.36
LlamaTouch W/O exact match 86.77 18.85 82.55 29.75 87.87 16.02
+ activity exact match 23.21 81.60 41.05 68.29 18.58 85.06
+ action exact match 86.62 19.01 81.81 30.49 87.87 16.02
+ UI component exact match 15.86 88.15 40.30 68.55 9.51 93.24
+ system state exact match 85.41 20.22 75.91 36.40 87.87 16.02
LlamaTouch W/O fuzzy match 10.54 93.26 23.36 84.02 7.21 95.66
+ screen-level fuzzy match 10.24 93.36 22.13 84.76 7.15 95.60
+ textbox fuzzy match 8.97 94.83 18.69 88.69 6.45 96.43
LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributing to the annotation and implementation of different essential state primitives. In this section, we evaluate the effectiveness of two types of match design in LlamaTouch: fuzzy match and exact match. We aggregate the execution traces of all mobile agents evaluated in $\S$§4.3. As for evaluation implementation, we disable exact match and fuzzy match separately, and then enable different primitives once per time. We present the results on (1) all tasks; (2) tasks in AITW (Rawles et al., 2023); and (3) tasks generated in LlamaTouch. The results of task completion rate and accuracy are shown in Table 6.
Exact match ensures key information in two screens are identical. As shown in the results, exact match contributes a significant portion of evaluation accuracy improvement. For example, without all exact match primitives, the evaluation accuracy of LlamaTouch on all tasks drops from 95% to 19%. Among all exact match primitives, activity match highly improves the validation accuracy. This is not only attributed to its simple yet efficient ability to locate functional application screens, but also the large portion of annotated activity primitives, i.e., 35% out of all primitives in our dataset. UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen. Action and system state exact match only improve a little accuracy as only a few tasks are annotated with these primitives.
Fuzzy match does not show notable improvement in evaluation accuracy, although there are 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset. Seldom of tasks with these primitives can be completed by the four agents we evaluated, leading to a slight improvement. Although primitives about fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world. With more performant mobile agents in the future, they will be further explored and evaluated.
4.5. Absolute Capabilities of Mobile Agents
In this section, we present the absolute capabilities of different mobile agents in mobile UI automation tasks. We first categorize tasks according to their sources: AITW and our self-constructed dataset. This aims to reveal whether an agent that has previously learned on a well-established dataset has the capability to adopt new applications/tasks. Second, we categorize tasks according to their difficulty, measured by the number of steps required to complete a task in the ground-truth dataset. Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 <<< steps $\leq$ 8), and high (step >>> 8).
Table 7. The comparison of task completion rate categorized by task sources (e.g., from the previous dataset AITW or from LlamaTouch’s generated dataset).
Agent End-to-end TCR
Overall AITW Generated
AutoUI 4.44 12.75 2.29
AutoDroid 14.84 22.77 12.79
AppAgent 10.91 21.57 8.14
CoCo-Agent 4.47 12.75 2.31
Performance among datasets. Among the four evaluated agents, AutoUI and CoCo-Agent are previously trained on AITW. AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively. We separately show end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to show their performance generalization to new scenarios (i.e., new tasks and new applications). Results in Table 7 show that AutoDroid and AppAgent also show higher TCR in AITW tasks than LlamaTouch’s generated tasks. This mainly attributes to the task complexity gap between the two datasets. AutoUI and CoCo-Agent achieve 12.75% TCR on tasks in AITW. However, for generated tasks in LlamaTouch, both of them operate poorly: with only 2.29% and 2.31% TCR, respectively. The TCR gap between datasets reveals their lack of capability to adapt to unseen tasks before. Considering massive real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
Performance under different task complexity. We categorize tasks into three difficulty levels according to the step number required to complete the task in the datasets. The results are shown in Table 8. Generally, mobile agents can better complete simple tasks that require fewer steps. There is a significant drop in TCR when task complexity increases. AutoDroid outperforms the other three agents for tasks under all difficulties. AppAgent achieves a similar TCR among all tasks as AutoDroid. We deem that the knowledgable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations. Therefore, agents based on GPT-4/4V achieve high TCR. However, there is still huge space for mobile agents to improve their capabilities in UI automation tasks.
Table 8. The comparison of task completion rate between mobile agents. Tasks are categorized to different difficulty levels according to the number of steps required to finished it in the ground-truth dataset.
Agent End-to-end TCR
[…]
AutoUI 4.44 4.95 4.19 4.82
AutoDroid 14.84 27.00 12.94 7.32
AppAgent 10.91 16.83 10.97 3.61
CoCo-Agent 4.47 7.92 3.91 2.41
5. Limitations of LlamaTouch
Supporting WebView-based apps. One major limitation of LlamaTouch is that its evaluation process requires Android’s view hierarchy to empower approximate UI layout matching and accurate UI component matching ($\S$§3.2). While some Android applications built with WebView (Google, 2024a) are unable to access their view hierarchy, making LlamaTouch cannot evaluate agent performance on these applications. However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing view hierarchy. When evaluating WebView-based apps are required, more advantaged techniques, such as screen similarity detection (Wu et al., 2023b, a), OCR/model-powered screen element recognition (Bunian et al., 2021; Xie et al., 2020), should be incorporated to retrofit LlamaTouch. We will consider this in future work.
Biases in identifying essential states. LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on pre-defined task execution traces. This process, however, could involve biases and leads to inaccuracy due to potentially limited knowledge in application execution and unseen circumstances. For example, given a task “Delete YouTube in the Google Play Store”, the pre-defined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution. However, they may ignore that this application might be not installed at all. In this case, there will be no explicit action of “clicking the uninstall button” while this task is still completed. Such biases might result in low accuracy in checking task completion rates. However, we think this limitation could be eliminated and refined by involving expert reviewing. Furthermore, for every single task description, involving diverse essential states on different potential task execution paths could better enhance the robustness of LlamaTouch’s evaluation design.
6. Conclusion
In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with faithfulness and scalability in UI automation tasks. LlamaTouch enables mobile agents to be tested on real mobile devices. At the evaluation stage, it only matches the task execution traces with annotated essential states. LlamaTouch well tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches. It achieves high evaluation accuracy like human validation, while preserving the scalability of evaluating on static datasets. By putting task execution on real mobile devices, we also reveal the limited capabilities of current mobile agents in UI automation tasks.
Appendix A Appendix
A.1. APIs Provided by LlamaTouch
LlamaTouch provides three categories of APIs for mobile agent integration and UI automation task execution. APIs, parameters, and their return values are listed in Table 9.
Table 9. APIs provided by LlamaTouch to mobile agents for UI automation task execution on real mobile devices.
API Category API Parameter Return Value
Metadata query get_task_instruction None str: A task description in natural language.
UI state query get_screenshot None
str: A base64 encoded string
representing the current screenshot.
get_view_hierarchy None
str: A string representing the
textual view hierarchy in XML.
Action space post_task_complete None None
post_task_impossible None None
post_press_home None None
post_press_back None None
post_click
x, y: Normalized x/y coordinates for a
targeted screen position.
None
post_type str: The text to be input. None
post_swipe
touch_x, touch_y: Normalized x/y
coordinates where the swipe begins.
lift_x, lift_y: Normalized x and y
coordinates where the swipe ends.
duration: The interval of the swipe gesture.
None

### v2

@@
-LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation
+LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Task Automation
 Abstract.
-The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in the task of mobile UI automation.
-However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined ones, are unscalable and unfaithful.
-To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device agent execution and faithful, scalable agent evaluation.
+The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in mobile UI task automation.
+However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined action sequences, are unscalable and unfaithful.
+To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device mobile UI task execution and faithful, scalable task evaluation.
 By observing that the task execution process only transfers UI states, LlamaTouch employs a novel evaluation approach that only assesses whether an agent traverses all manually annotated, essential application/system states.
-LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with real mobile environments for task completion.
+LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with realistic mobile environments for task execution.
 (2) Fine-grained UI component annotation that merges pixel-level screenshots and textual screen hierarchies to explicitly identify and precisely annotate essential UI components with a rich set of designed annotation primitives.
-(3) A multi-level state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen with unpredictable UI layout/content dynamics.
-LlamaTouch currently incorporates four mobile agents and 495 UI automation tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones for more diverse mobile applications.
-Evaluation results demonstrate the LlamaTouch’s high faithfulness of evaluation in real environments and its better scalability than human validation.
+(3) A multi-level application state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen, even with unpredictable UI layout/content dynamics.
+LlamaTouch currently incorporates four mobile agents and 496 tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones to cover more diverse mobile applications.
+Evaluation results demonstrate LlamaTouch’s high faithfulness of evaluation in real-world mobile environments and its better scalability than human validation.
 LlamaTouch also enables easy task annotation and integration of new mobile agents.
 Code and dataset are publicly available at https://github.com/LlamaTouch/LlamaTouch.
-copyright: none
+mobile agent, UI task automation, evaluation, testbed
+journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676382isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Human computer interaction (HCI)
 1. Introduction
-Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious, cumbersome smartphone operations.
+Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious and cumbersome smartphone operations.
 These agents are particularly beneficial for individuals with visual or hand impairments, or in situations where using a screen is not practical (e.g., driving).
 Notable mobile agents, such as Apple Siri (Apple, 2024) and Google Assistant (Google, 2024b), have become indispensable services on smartphones.
-The recent advent of large language models (LLMs) and multimodal LLMs have facilitated researchers building more powerful mobile agents (Li et al., 2024; Wen et al., 2023; Wang et al., 2023; Lee et al., 2023).
-The key capability of those agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI automation, e.g., “forward the last email from Bob to Alice”.
-Despite claims of powerful UI automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed.
-Unlike traditional machine learning models evaluated on well-established, static datasets, mobile agents need to interact with the dynamic, indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs.
-Simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
+The recent advent of large language models (LLMs) and multimodal LLMs has facilitated researchers in building more powerful mobile agents (Li et al., 2024; Wen et al., 2024; Wang et al., 2023; Lee et al., 2023).
+The key capability of these agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI task automation, e.g., “forward the last email from Bob to Alice”.
+Despite claims of powerful task automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed.
+Unlike traditional machine learning models evaluated on well-established static datasets, mobile agents need to interact with the dynamic and indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs.
+Additionally, mobile devices use touch and gesture-based interactions (e.g., swipe, pinch), leading to diverse and ambiguous inputs.
+This variability complicates static action-based matching algorithms.
+Therefore, simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
 In general, there are two methods to evaluate mobile UI automation tasks, but neither achieves both high faithfulness and scalability.
 (1) The most intuitive approach is to request humans to verify the completion of tasks.
-However, human evaluation is difficult to reproduce (Chiang and Lee, 2023).
-The requisite human effort increases with the number of agents, tasks, and evaluation platforms.
-(2) The most popular approach used in most prior arts (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations.
+However, human evaluation is difficult to reproduce (Chiang and Lee, 2023), and the requisite human effort increases with the number of agents, tasks, and evaluation platforms.
+(2) The most popular approach used in most prior work (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2024; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations.
 The key idea is to ask annotators to generate a correct sequence of actions that succeed on the task as data labels, and then compare agent-generated actions to these labels.
@@
 Consequently, it leads to a significantly higher false negative rate.
-For example, for task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged.
+For example, for the task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged.
 Only taking one of these execution paths as the reference may incorrectly verify a task that is essentially completed.
-Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI automation capabilities, yet is impossible to evaluate in a static dataset.
-These limitations are further demonstrated in $\S$§$2.$
-This paper presents LlamaTouch, the first testbed for evaluating mobile agents on real-world mobile environments without compromising faithfulness and scalability.
+Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI task automation capabilities, yet is impossible to evaluate in a static dataset.
+These limitations are further demonstrated in §$\S$2.2.
+This paper presents LlamaTouch, the first testbed for evaluating mobile agents in real-world mobile environments without compromising faithfulness and scalability.
 The key idea of LlamaTouch is to check the task execution trace against a few “essential states” identified by the annotators, rather than matching them against predefined action sequences in static traces.
-For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened; and (2) the application is on the login page.
+For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened, and (2) the application is on the login page.
 Other operations, like app installation, are considered non-essential and should be ignored.
-During task execution, LlamaTouch enables mobile agents to retrieve only the task description from the static dataset, while device states are directly acquired from real-world devices.
-Actions produced by mobile agents are executed on those smartphones, and all UI interaction data are recorded as task execution traces.
+During task execution, LlamaTouch enables mobile agents to retrieve only task descriptions from static datasets, while device states are directly acquired from realistic mobile devices.
+Actions produced by mobile agents are directly operated on those devices, and all UI interaction data are recorded as task execution traces.
 In the evaluation phase, LlamaTouch compares task execution traces with annotated essential states to determine whether a task has been completed.
@@
 It combines pixel-level screenshots and textual screen hierarchies to explicitly highlight important UI components.
-With a rich set of annotation primitives provided by LlamaTouch, it reduces human efforts to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the content of a textbox should be exactly matched.
-These annotated primitives are subsequently used for faithful evaluation.
-(2) During the evaluation process, LlamaTouch employs a multi-level state match algorithm that combines fuzzy and exact matches on diverse annotated UI states.
-It uses (i) approximate screen match that enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state match that detects and matches critical on-screen information.
+With a rich set of annotation primitives provided by LlamaTouch, it reduces human effort to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the text inside a textbox should be exactly matched.
+These annotated UI states are subsequently used for faithful evaluation.
+(2) During evaluation, LlamaTouch employs a multi-level state matching algorithm that combines fuzzy and exact matches on diverse annotated UI states.
+It uses (i) approximate screen matching, which enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state matching, which detects and matches critical on-screen information.
 Dataset and testbed.
-We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world Android environments.
-This dataset includes 495 distinct tasks encompassing a wide array of popular Android applications.
-We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with real Android environments.
+We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world mobile environments.
+This dataset includes 496 distinct tasks encompassing a wide array of popular Android applications.
+We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with realistic Android environments.
 This testbed provides a collection of concise, widely used APIs, ensuring compatibility with most mobile agents.
-Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI automation tasks in real-world scenarios.
+Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI task automation in real-world scenarios.
 Evaluation.
-We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) as actual smartphones.
-Currently, LlamaTouch has four built-in agents, AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2023), and CoCo-Agent (Ma et al., 2024), along with 495 diverse tasks.
-The number of supported agents and tasks will be extended in the future.
-With human validation as the ground truth for task completion, LlamaTouch achieves nearly 80% accuracy in detecting completed tasks in real-world environments, while exact action match on static datasets fails to do so.
+We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) and one Google Pixel 5 smartphone as realistic Android environments.
+Currently, LlamaTouch has four built-in agents, including AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2024), and CoCo-Agent (Ma et al., 2024), along with 496 diverse tasks.
+With human validation results as the ground truth for task completion, LlamaTouch achieves nearly 80% evaluation accuracy in detecting completed tasks in real-world environments, while prior action-based evaluation methods fail to do so.
 We also reveal the limitations of current mobile agents in handling tasks practically in real-world environments.
 Contributions are summarized as follows.
-  • We observed the weakness of high false negative rates in evaluating mobile UI automation agents using static datasets.
+  • We observed the weakness of high false negative rates in evaluating mobile UI task automation agents using static datasets.
 To address this, we proposed an evaluation design that only compares essential states rather than concrete action sequences.
   • We devised a method for annotating essential states using a variety of annotation primitives.
-This approach combines visually intuitive screenshots with semantically precise VHs to enable fine-grained and accurate UI component localization and annotation.
-  • We designed a novel task evaluation system that employs both exact and fuzzy matching at various UI state levels.
-It enables faithful evaluation of mobile agents and scales well to dynamic execution environments.
-  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI automation agents in real-world mobile environments.
-It comprises 495 UI automation tasks with human-annotated essential states.
+This approach combines visually intuitive screenshots with semantically precise view hierarchies to enable fine-grained and accurate UI component localization and annotation.
+  • We designed a novel task evaluation approach that employs both exact and fuzzy matching at various UI state levels.
+It enables faithful evaluation of mobile agents and adapts well to dynamic execution environments.
+  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI task automation agents in real-world mobile environments.
+It comprises 496 tasks with human-annotated essential states.
 Four agents integrated in LlamaTouch demonstrate its faithfulness and scalability in UI automation task evaluation.
-2. Background and Motivation
-2.1. Agents for Mobile UI Automation
-Mobile agents have simplified the cumbersome and dull operations on smartphones for users.
-The progression of mobile agents for mobile UI task automation can be categorized into three phases.
-(1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through pre-defined application programming interfaces.
-This approach is reliable while limited in structured and predictable tasks.
-(2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities were still confined by their training data.
-(3) Recently, LLMs and multi-modal LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2023; Wang et al., 2023).
-These models, owing to their vast knowledge base, can understand complex, real-world scenarios.
-Mobile agents powered by these models are able to interpret natural language instructions accurately and translate them into actionable tasks on smartphones’ screens.
-This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
-Mobile agents typically operate with the following components.
-Controller is the brain of mobile agents.
-It interprets task instructions and UI contexts, and generates actions to be executed on the current UI context.
-The widely-used controllers include DL models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (GPT-3.5, GPT-4) (Wen et al., 2023; Wang et al., 2023; Lee et al., 2023), and multi-modal LLMs (GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
-Input: UI representation.
-Existing mobile agents take a task description and a UI representation as the input of their controller.
-There are two basic types of UI representations: screenshots and VH.
-A screenshot is a visual capture of the current screen.
-A text-format VH provides a tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents.
-On top of screenshots and VH, some controllers further extract UI semantics to enhance UI understanding.
-For example, Yan et al.
-(Yan et al., 2023) overlies numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) also converts screenshots to bounding boxes and labels, enabling them comprehensible by LLMs.
-Further processing based on VH, like converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2023).
-Output: actions.
-The output of controllers are actions to be executed on the current screen, such as click, swipe, and input text.
-Action parameters can be abstracted at different levels depending on the agent’s design and their input format.
-(1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023).
-It operates as a direct interaction with the screen like human operations.
-(2) Icon marker (Yan et al., 2023; Yang et al., 2023).
-The output target will specify a specific icon or graphical element within the UI representation.
-(3) HTML index (Wang et al., 2023; Wen et al., 2023; Lee et al., 2023).
-By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
 Table 1.
 The comparison between mobile agent benchmarks.
-LlamaTouch is the first testbed designed for mobile agents driven by essential state match.
-LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of diverse UI states.
+LlamaTouch is the first testbed designed for mobile agents driven by essential state matching.
+LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of matching implementations.
 Benchmark Platform
@@
 LlamaTouch Mobile
-Figure 1. Two major limitations of evaluating mobile agents on static datasets.
-$2.$ Mobile UI Automation Benchmarks
-As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI automation tasks, while none of them achieve both faithfulness and scalability.
-Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets (task descriptions and UI representations).
-Mobile agents predict concrete actions on static UI representations, which will be compared with the ground-truth actions in datasets.
+2. Background and Motivation
+2.1. Agents for Mobile UI Task Automation
+Mobile agents have simplified the cumbersome and dull operations on smartphones for users.
+The progression of mobile agents for mobile UI task automation can be categorized into three phases.
+(1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through predefined application programming interfaces.
+This approach is reliable while limited in structured and predictable tasks.
+(2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities are still confined by their training data.
+(3) Recently, LLMs and multi-modality LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2024; Wang et al., 2023).
+These models, owing to their vast knowledge base, can understand complex, real-world mobile screens.
+Mobile agents powered by these models can accurately interpret natural language instructions and translate them into actionable tasks on smartphone screens.
+This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
+Mobile UI task automation agents typically operate with the following components.
+Controller is the brain of mobile agents.
+It interprets task instructions and UI contexts, and then generates actions to be executed on the current UI context.
+Widely-used controllers include deep learning models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (e.g., GPT-4, Llama) (Wen et al., 2024; Wang et al., 2023; Lee et al., 2023), and multi-modality LLMs (e.g., GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
+Input: UI Representation.
+Existing mobile agents take a task description and UI representations as the input of their controller.
+There are two basic types of UI representations: screenshot and view hierarchy (VH).
+A screenshot is a visual capture of the current screen.
+A VH provides a textual tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents.
+On top of screenshots and VHs, some controllers further extract UI semantics to enhance UI understanding.
+For example, Yan et al.
+(Yan et al., 2023) overlay numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) converts screenshots to bounding boxes and labels, making them comprehensible to LLMs.
+Further processing based on VH, such as converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2024).
+Output: Action.
+The output of controllers consists of actions to be executed on the current screen, such as click, swipe, and input text.
+Action parameters can be abstracted at different levels depending on the agent’s design and input format.
+(1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023): This operates as a direct interaction with the screen, similar to human operations.
+(2) Icon marker (Yan et al., 2023; Yang et al., 2023): The output target will specify a specific icon or graphical element within the UI representation.
+(3) HTML index (Wang et al., 2023; Wen et al., 2024; Lee et al., 2023): By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
+(a) Inaccurate action match in two tasks. Failed reasons: “Left”: wrong click parameter; “Right”: wrong input text.
+(b) Different task execution paths lead to the same screen.
+Figure 1. Two major limitations of evaluating mobile UI task automation on static datasets.
+2.2. Mobile UI Task Automation Benchmarks
+Figure 2.
+LlamaTouch workflow and code demonstrations for mobile UI task execution and trace evaluation.
+LlamaTouch enables mobile UI task automation agents to integrate easily with AgentEnv for on-device task execution with minimal programming effort.
+Agent execution traces recorded by AgentEnv are used in conjunction with the LlamaTouch dataset in a separate evaluation process.
+As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI task automation, but none of them achieve both faithfulness and scalability.
+Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets with task descriptions, UI representations, and actions sequences.
+Mobile agents predict concrete actions on static UI representations, which are compared with the ground-truth actions in the datasets.
 While this approach is straightforward, it is insufficient to reveal the performance of mobile agents for two reasons.
-(1) Inaccurate exact action match are pervasive when a functionally correct action is deemed incorrect due to different action parameters.
-Figure 1(a) shows two cases.
-First, for click actions, the clickable area defined in the dataset may be narrow.
-While in reality, the area might encompass the entire UI component (marked with the red bounding box).
-Second, non-identical text inputs can lead to the same correct result (marked with green bounding boxes) in most search tasks.
-This issue has also been observed in previous literature but has not been solved (Yan et al., 2023; Rawles et al., 2023).
+(1) Inaccurate exact action match.
+Functionally correct actions may be deemed incorrect due to different action parameters.
+Figure 1 illustrates two cases.
+First, for click actions, the clickable area defined in the dataset may be narrow, whereas in reality, the area might encompass the entire UI component (marked with the red bounding box).
+Second, non-identical text inputs can lead to the same correct search result (marked with green bounding boxes) in most search tasks.
+This issue has been observed in previous literature but remains unsolved (Yan et al., 2023; Rawles et al., 2023).
 (2) Lack of tolerance for different execution paths.
-In real-world environments, an UI automation task can usually be completed in various paths, as demonstrated in Figure 1(b1) and Figure 1(b2).
-However, pre-defined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
-There are also work that enable agent execution in real-world environments.
-AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023) enable the execution of mobile automation tasks in real environments.
-However, they do not essentially support essential state match during the end-to-end task execution, therefore compromising evaluation accuracy.
-AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets does not fully tolerant redundant actions in task execution paths.
-They proposed a subsequence-based action match: a task is treated as completed when it contains the ground-truth action sequence as its subsequence.
-We take AndroidArena as a baseline in $\S$§4.3 to compare its evaluation accuracy with LlamaTouch.
-WebArena (Zhou et al., 2023) provides a real playground for Web agents.
+In real-world environments, a task can usually be completed in various paths based on different device/application states, as shown in Figure 1.
+However, predefined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
+There is also other work that enables agent execution in real-world environments, such as AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023).
+However, they do not inherently support essential state match during end-to-end task execution, therefore compromising evaluation accuracy.
+AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets: it does not fully tolerate redundant actions in task execution paths.
+They proposed a subsequence-based action match, where a task is treated as completed if it contains the ground-truth action sequence as its subsequence.
+We take AndroidArena as a baseline in §$\S$5.3 to compare its evaluation accuracy with LlamaTouch.
+WebArena (Zhou et al., 2023) provides a realistic playground for web agents.
 It uses essential states to evaluate task completion (e.g., the final result should be or should include some key information).
-LlamaTouch differs from WebArena on the mobile platform in both the essential state annotation and evaluation process: (i) LlamaTouch combines the visual screenshot with textual view hierarchy of the same screen for fine-grained and precise UI component identification; (ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
+LlamaTouch differs from WebArena on the mobile platform in both essential state annotation and evaluation process: (i) LlamaTouch combines visual screenshots with textual VHs of the same screens for fine-grained and precise UI component identification.
+(ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
 Human validation is usually used to validate whether a UI automation task is completed (Rawles et al., 2023; Taeb et al., 2023).
-However, the cost of human validation is too high, thereby scaling poorly to multiple tasks, agents, and mobile devices.
-LlamaTouch ensures high scalability as evaluating on static datasets while preserving faithfulness as human validation.
-Figure 2. The workflow of on-device task execution.
+However, the cost of human validation is too high, scaling poorly to multiple tasks, agents, and mobile devices.
+LlamaTouch ensures high scalability as with evaluating on static datasets while preserving faithfulness similar to human validation.
 3. LlamaTouch Design
+Figure 2 shows the workflow of using LlamaTouch for on-device task execution and trace evaluation, using the well-constructed LlamaTouch dataset.
 Compared to previous evaluation approaches, LlamaTouch exhibits the following benefits.
-$\bullet$ Practical on-device task execution ($\S$§3.1).
-Most of the previous evaluation methods are simulated on mobile UI interaction datasets.
-LlamaTouch, on the other side, enables mobile agent evaluation on real mobile devices, revealing their real capabilities in the wild.
-$\bullet$ Fine-grained essential application state annotation ($\S$§3.2).
-By observing UI automation task execution only transfers essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched during task execution.
-$\bullet$ Faithful and scalable task evaluation ($\S$§3.3).
-LlamaTouch evaluates mobile agents by comparing task execution traces with annotated application states.
-By combining exact matching with fuzzy matching algorithms on different application states, LlamaTouch achieves faithful and scalable task evaluation.
-3.1. On-device Task Execution
-Evaluating mobile agents on pre-defined traces, as discussed in Section 2.2, depicts significant inaccuracy.
-LlamaTouch empowers on-device UI automation task execution to reveal the real capabilities of mobile agents.
-To simplify this process, LlamaTouch operates as a bridge between existing mobile agents and real mobile devices (e.g., smartphones, Android emulators, and cloud device farms).
-As shown in Figure 2, mobile agents execute tasks on LlamaTouch by taking the following processes.
-1 Mobile agents request one task instruction from LlamaTouch.
-The instruction comes from one supported dataset by LlamaTouch, e.g., AITW.
-2 With this task instruction, mobile agents will then request UI representation (e.g., screenshot, VH) from LlamaTouch, which will be 3 forwarded to mobile devices.
-4 Taking the task instruction and UI representation as inputs, mobile agents will predict an action to be performed on the current UI, and dispatch predicted actions to LlamaTouch.
-5 LlamaTouch forwards and executes the agent-predicted action to real mobile devices.
-The processes from 2 to 5 are repeatedly executed until mobile agents treat the task as completed.
-During task execution, all UI representations and corresponding actions are captured for further evaluation in $\S$§$3.$ These UI representations (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application at the foreground of each screen; and (4) actions performed on every screen.
-Essential system states such as the list of installed applications will also be dumped for faithful mobile agent evaluation, as we will illustrate in $\S$§3.2.
-Appendix A.1 details APIs provided by LlamaTouch for mobile agent integration.
-Table 2. Primitives of essential states in LlamaTouch that are used in the annotation and evaluation processes.
+$\bullet$ Practical on-device mobile UI task execution (§$\S$3.1).
+Most previous evaluation methods are simulated on mobile UI interaction datasets.
+LlamaTouch, on the other hand, enables mobile agents to operate in realistic mobile environments for UI task automation, revealing their true capabilities in real-world scenarios.
+$\bullet$ Fine-grained essential application state annotation (§$\S$3.2).
+By observing UI automation task execution transfers states of essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched for task completion.
+This approach avoids the previous static evaluation methods’ focus on the determinism of traces, reducing the probability of false negatives in evaluation.
+$\bullet$ Faithful and scalable task evaluation (§$\S$3.3).
+LlamaTouch evaluates the performance of mobile agents by comparing their task execution traces captured in real-world mobile environments with annotated essential states.
+By combining exact and fuzzy matching algorithms on different application states, LlamaTouch achieves faithful task evaluation without losing scalability.
+3.1. On-device Mobile UI Task Execution
+Table 2. A set of primitives used in essential state annotation and trace evaluation.
 Match Type State Type Primitive Keyword Use Case
-Fuzzy match UI state Screen info fuzzy<-1> Check if the contents on two screens are approximately identical.
+Fuzzy match UI state Screen info fuzzy<-1> Verify if the contents on two screens are approximately identical.
 Textbox fuzzy<n>
-Check if the content of the target textbox is semantically similar
+Verify if the content of the target textbox is semantically similar
 to the content of the original textbox<n> in the ground-truth UI.
 Exact match Activity activity
-A coarse-grained approach to determining if two UIs represent
+A coarse-grained approach to determine whether two UIs represent
 the same functional screen in an application.
@@
 exclude<n>
-Check if the UI component is exactly identical to the UI component<n>,
-or does not occur, in the ground-truth UI.
+Verify if the UI component is exactly identical to the UI component<n>,
+or if it does not occur, in the ground-truth UI.
 System state (Un)installation
@@
 uninstalled<app>
-Check if the target application named "app" has been successfully
-installed/uninstalled.
+Verify if the target application named "app" has been successfully
+installed or uninstalled.
 Action Action
@@
 type<input_text>
-Check if two actions and their parameters are identical.
-Figure 3.
-Annotated essential state for task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes that should be exactly matched.
-The essential state represents the application state after task execution: “the shopping cart on bestbuy is empty”.
+Verify if two actions and their parameters are identical.
+Evaluating mobile agents on predefined, deterministic traces, as discussed in Section 2.2, depicts significant inaccuracy.
+LlamaTouch empowers on-device mobile UI task execution to reveal the real capabilities of mobile agents.
+To simplify this process, we propose AgentEnv, a bridge between existing mobile agents and realistic mobile environments (e.g., smartphones, Android emulators, and cloud device farms).
+With AgentEnv, mobile agents execute tasks on real-world mobile environments by following the processes shown in Figure 3.
+1 Mobile agents request a task instruction from AgentEnv.
+The instruction comes from the LlamaTouch dataset.
+2 With the task instruction, mobile agents request UI representations (e.g., screenshot, VH) from AgentEnv, which are then 3 forwarded to mobile devices.
+4 Taking the task instruction and UI representations as inputs, mobile agents predict an action to be performed on the current UI and dispatch the predicted action to AgentEnv.
+5 AgentEnv forwards and executes the agent-predicted action to mobile environments.
+The processes from 2 to 5 are repeated until mobile agents consider the task completed.
+During task execution, all UI representations and corresponding actions are captured as task execution traces for further evaluation in §$\S$3.3.
+These UI representations include (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application in the foreground of each screen; and (4) actions performed on each screen.
+Essential system states, such as the list of installed applications, will also be recorded for faithful mobile agent evaluation.
+The left-hand side of Figure 2 demonstrates the code implementation of the interaction between mobile agents and AgentEnv.
+Appendix A.1 presents details of the APIs provided by LlamaTouch for mobile agent integration.
+Figure 3. Interaction between mobile agents, AgentEnv, and real-world mobile environment in the on-device mobile UI task execution process.
 3.2. Essential Application State Annotation
-Two major cases demonstrated in $\S$§2.2 make exact matching on pre-defined traces unfaithful.
+Two major cases demonstrated in §$\S$2.2 highlight why exact action match on predefined action sequences is unfaithful.
 First, it demands that the agent-generated actions and their parameters match exactly with those in the dataset.
-Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, therefore unable to evaluate alternative task execution paths.
+Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, making it unable to evaluate alternative task execution paths.
 Insight: The task execution process transfers identifiable application states.
-The internal of the execution process of a UI automation task is changing application states that some of them can be explicitly represented by UI components.
+During the execution of a UI automation task, the application states change, and some of these states can be explicitly represented by UI components.
 Even if the task execution paths differ, there are overlaps in the essential application states.
 We can utilize these overlapping application states to determine whether a task achieves some milestones or is completed.
-As the example shown in Figure 1(b), by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page.
+As the example shown in Figure 1, by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page.
 Other actions, like detecting whether the application is installed, can be omitted during evaluation.
-To do so, essential states should be accurately identified and annotated.
-However, simply annotating application states at the whole UI representation level (e.g., a screenshot) and comparing screen-level similarity (Feiz et al., 2022), are far coarse-grained and may lead to inaccurate matching.
-For example, the screen contents of a web-shopping application could be subtly different due to slight movement or dynamically loaded contents across different executions.
-For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation (i.e., screenshot) to separate UI components.
-LlamaTouch simplifies the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components.
+To achieve this, essential states should be accurately identified and annotated.
+However, simply annotating application states at the whole UI representation level (e.g., an entire screenshot) and comparing screen-level similarity (Feiz et al., 2022) is too coarse-grained and may lead to inaccurate matching.
+For example, the screen contents of a web-shopping application could be subtly different due to nondeterministic swiping gestures or dynamically loaded contents across different executions.
+For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation into separate UI components.
+This is achieved by simplifying the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components.
 These extracted UI components are combined with visually intuitive screenshots to provide precise overlayed bounding boxes and unique identifiers to annotators.
-Figure 3 shows an example of the VH-enhanced screenshot provided to annotators.
-Primitives.
+Figure 4 shows an example of VH-enhanced screenshots in a task provided to annotators.
+Annotation primitives.
 LlamaTouch incorporates a list of primitives for essential state annotation.
-These primitives are used to extract essential information on the screen.
-Annotators are instructed to clearly identify and use primitives to represent application states that are informative and deterministic to validate the task execution results.
+These primitives represent essential information on the screen and indicate how this information should be matched.
+Annotators are responsible for clearly identifying and using primitives to represent application states that are informative and deterministic for validating task execution results.
 Table 2 comprehensively shows these primitives and their use cases.
-Overall, LlamaTouch incorporates six types of primitives and nine keywords for annotation.
-These primitives can be divided into three types according to the state type they represent.
+Currently, LlamaTouch incorporates six types of primitives and nine keywords for annotation.
+These primitives can be divided into three types according to the application state type they represent.
+Figure 4.
+Annotated essential states for the task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes with the exact keyword.
+The essential states represent the application state after task execution: “the shopping cart on bestbuy is empty”.
 $\bullet$ UI state.
 The purpose of UI state annotation is to extract and compare whether two screens contain identical or similar information.
-As a screen may contain different types of components, LlamaTouch has four major primitives to cover all these UI components.
+As a screen may contain different types of components, LlamaTouch uses four major primitives to cover the comparison logic of all these UI components.
 (1) Screen info and (2) textbox are used to compare whether the whole screen or a dedicated textbox has similar contents.
@@
 (4) UI component is used to deterministically annotate the state of a UI component, including textbox, button, image, etc.
-For example, if the content of a textbox should be identical, it should be annotated with the exact keyword.
-States that require certain UI components should not occur on the screen should be annotated with the exclude keyword.
+For example, if the content of a textbox should be matched, this textbox should be annotated with the exact keyword.
+States that require certain UI components not to occur on the screen can be annotated with the exclude keyword.
 $\bullet$ System state.
-LlamaTouch involves system state annotation mainly inspired by the tasks in a previously built dataset: AITW (Rawles et al., 2023).
-For example, a task “install app YouTube Kids” can be detected directly using shell commands (e.g., pm list packages), without involving redundant UI operations.
-Currently, LlamaTouch supports programmatically checking application installation.
-The keywords can be easily extended when other system states are required to be detected.
+System state annotation in LlamaTouch is primarily inspired by tasks from the previously built dataset: AITW (Rawles et al., 2023).
+For example, a task like “install app YouTube Kids” can be detected directly using shell commands to access system states (e.g., pm list packages), without involving complex UI states.
+Currently, LlamaTouch supports programmatically checking application installation status.
+The keywords can be easily extended and customized to detect other system or application states.
 $\bullet$ Action.
 Although the principle of annotation is to detect application state transfer during task execution.
-Sometimes, action on a specific UI representation is necessary to validate agent behaviour, especially when there is not enough state identifier shown in UI.
-LlamaTouch provides click and type keywords that covers most common actions.
+Sometimes, action on a specific UI representation is necessary to validate agent behavior, especially when there are not enough state identifiers shown in the UI.
+LlamaTouch provides click and type keywords that cover the most common actions.
 Case study.
 Essential state annotation can be divided into the following processes.
-First, for each UI representation, LlamaTouch overlayed all functional UI components with numeric markers to the screenshot as shown in Figure 3.
+First, for each UI representation, LlamaTouch overlays all functional UI components with numeric markers on the screenshot as shown in Figure 4.
 This is done by extracting the precise metadata of each UI component from textual VH.
-Then, annotators will identify what are essential states that should be checked during evaluation to ensure the task is completed.
-LlamaTouch simplify this process by only requiring annotators to explicitly identify what UI element should be matched.
+Annotators will then identify the essential states that should be checked during evaluation to ensure the task is completed.
+LlamaTouch simplifies this process by only requiring annotators to explicitly identify what UI element should be matched and what annotation primitive should be used.
 For example, after emptying the shopping cart on bestbuy, the screen will display a textbox with the content “Your cart is empty”.
@@
 The keyword exact<27> is used to validate whether two screens are both in the shopping cart of bestbuy.
-All annotations along with the UI representations and task descriptions construct the essential state-powered dataset.
+All annotations, along with the UI representations and task descriptions, construct the essential state-powered dataset.
 These essential states are used by LlamaTouch for faithful task evaluation.
-$3.$ Faithful and Scalable Task Evaluation
-Annotated essential states ($\S$§3.2) and task execution traces captured on real devices ($\S$§3.1) are used for evaluation on LlamaTouch.
-LlamaTouch iterates essential states to determine whether a task execution trace goes through all annotated states sequentially.
-If so, this task is deemed completed.
+3.3. Faithful and Scalable Task Evaluation
+Annotated essential states (§$\S$3.2) and captured task execution traces (§$\S$3.1) are jointly used for evaluation in LlamaTouch.
+LlamaTouch iterates through essential states to determine whether a task execution trace sequentially matches all annotated states.
+If so, the task is deemed completed.
 To achieve faithful evaluation, the most significant challenge lies in how to ensure the task execution trace matches the essential states.
-Task execution traces captured from real-world environments may contain dynamic screen content; it is vital to adapt the annotated essential states on a static dataset to dynamic, real-world contents, while achieving precise match on only critical information.
-To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components, to ensure faithful evaluation.
-It first approximately matches two screens according to their UI representation and activity.
-Then, within two matched screens, it compares the annotated primitives that should be fuzzily or exactly matched.
+Task execution traces captured from real-world mobile environments may contain dynamic screen content; it is vital to adapt the annotated essential states from a static dataset to varying screens captured from the real world, while achieving precise matches on only critical information.
+To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components to ensure faithful evaluation.
+The algorithm first approximately matches two screens according to their UI representations and activities.
+Then, within two matched screens, it compares each annotated essential state with iterated UI components in the target screen based on their annotated primitives.
+A task is considered completed only when all annotated essential states have matching counterparts.
+The matching philosophy of all annotation primitives is as follows.
 Approximate screen match is used to ensure two screens are on the same functionality page of an application, even under high screen content dynamics.
-This is the prior process for in-screen exact/fuzzy content match.
-To achieve this, LlamaTouch utilizes two annotation primitives proposed in $\S$§3.2: activity and screen info.
+This is the preliminary process for in-screen exact/fuzzy content match.
+To achieve this, LlamaTouch utilizes two annotation primitives proposed in §$\S$3.2: activity and screen info.
 1.
@@
 Activity represents an entry point for users to interact with the application (and, 2024).
-This is an obvious identifier to indicate whether two screens are in the same application.
+This is a clear identifier to indicate whether two screens are in the same application.
 Screens with distinct functionalities typically have unique activity names, even if they are within the same application.
-For example, the activity of the main setting page is com.android.settings.Settings, while the value is com.wifiadmin.settings.WifiSettingsActivity in the Wi-Fi setting page.
+For example, the activity of the main settings page is “com.android.settings.Settings”, while the Wi-Fi settings page has the activity “com.wifiadmin.settings.
+WifiSettingsActivity”.
 Typically, exact application activity match acts as a foundational filtering process to identify whether two screens are on the same page.
-However, some specific application design philosophies may make different functional pages of an application contain the same activity name (sin, 2018).
-Under this circumstance, LlamaTouch involve screen info for accurate screen match.
+However, some specific application design philosophies may cause different functional pages of an application to contain the same activity name (sin, 2018).
+Under these circumstances, LlamaTouch involves screen info for accurate screen match.
 2.
-Fuzzy screen info match is utilized when the activity name cannot reveal different functional pages of an application.
-To better compare two screens, the primary thing is to only extract critical information from the screen.
-In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation as in prior work (Wen et al., 2023; Wang et al., 2023), while containing the types of every single UI component.
-LlamaTouch compares two simplified HTML representations of two screens with a cosine similarity of their sentence embeddings.
-If the result falls within a threshold, LlamaTouch treats two screens are matched.
-Two screens are similar when they are above a predefined threshold, e.g., 0.85 in our experiments.
+Fuzzy screen info match is utilized when the activity name cannot differentiate between functional pages of an application.
+To better compare two screens, it is crucial to extract only the critical information from the screen.
+In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation, as in prior work (Wen et al., 2024; Wang et al., 2023), while preserving the types of every UI component.
+LlamaTouch compares two simplified HTML representations of the screens using the cosine similarity of their sentence embeddings.
+Two screens are deemed similar when their cosine similarity exceeds a predefined threshold, e.g., 0.85 in our experiments.
 The fuzzy screen info match design helps LlamaTouch maintain faithfulness during evaluation when dealing with dynamic screen contents.
+Figure 5.
+Workflow for constructing the LlamaTouch dataset.
+Generally, the workflow has three independent processes.
+(1) Generate task descriptions by sampling from previous datasets or constructing new ones through humans or LLMs.
+(2) Record UI interaction traces according to task descriptions.
+(3) Annotate essential states atop UI interaction traces.
 Mixed UI state match will be applied to matched screens after the approximate screen match process.
@@
 1.
-Exact UI component match requires the annotated UI component should be identical in two screens to be compared, including all their attributes such as class, text, and selected in VH.
+Exact UI component match requires that an annotated UI component be identical in two screens being compared, including all their attributes such as class, text, and selected in VH.
 This is especially useful for evaluating the content of a textbox, the status of a button (e.g., checked or not), or a selectable icon on the screen.
-Given a target UI component to be exactly matched, LlamaTouch will iterate nodes in VH of the matched screen, until one matched UI component is found.
-Exact UI component match fails if there is no matched node on available screens.
-In our dataset, exact UI component match contributes to 50.5% (688 out of 1,362) of annotated essential states.
+Given a target UI component to be exactly matched, LlamaTouch will iterate through nodes in the VH of the matched screen until a matching UI component is found.
+Exact UI component match fails if no matching node is found on available screens.
+In our dataset, exact UI component match accounts for 51% (698 out of 1,379) of annotated essential states.
 2.
-Fuzzy textbox match is crucial for comparing content inside a textbox on the screen, especially when the content may be slightly different.
-Semantically similar searching keywords with the same intent that complies with a specific task description should be matched, as they will lead to the same results.
-For example, Figure 1(b) shows searching “Microsoft Excel” and “Excel” in the Google Play Store both display the target application.
-LlamaTouch extracts the content of the annotated textbox, and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
+Fuzzy textbox match is crucial for comparing the content inside a textbox on the screen, especially when the content may be slightly different.
+Semantically similar search keywords with the same intent that comply with a specific task description should be matched, as they will lead to the same results.
+For example, Figure 1 shows searching for “Microsoft Excel” and “Excel” in the Google Play Store both display the target application.
+LlamaTouch extracts the content of the annotated textbox and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
 3.
 Action match.
-Although the initiative of LlamaTouch is to detect state transfer during task execution.
-There are still cases for evaluating concrete actions performed on the screen, such as clicking one specific UI component, or typing the correct captcha.
-LlamaTouch directly compares actions and their parameters of the annotated ones with their counterparts on the matched screens.
+Although the initiative of LlamaTouch is to detect state transfer during task execution, there are still cases for evaluating concrete actions performed on the screen, such as clicking a specific UI component or typing the correct captcha.
+LlamaTouch directly compares the actions and their parameters of the annotated actions with their counterparts on the matched screens.
+Different from prior studies (Rawles et al., 2023), LlamaTouch uses XPath (xpa, 2017) of target UI components extracted from screen VHs as parameters for click actions.
+This approach provides more tolerance in the action space compared to using precise coordinates.
 4.
@@
 LlamaTouch currently supports checking whether an application is installed or not.
-Such system state is recorded during task execution in real-world mobile devices ($\S$§3.1).
-LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace.
+Such system states are recorded during task execution on real-world mobile devices (§$\S$3.1).
+LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace, which is also recorded by AgentEnv.
 Through the multi-level state matching algorithm, LlamaTouch achieves high evaluation accuracy on real-world task execution traces, while preserving the scalability of evaluating on static datasets.
-Experiments in $\S$§4 show the faithful evaluation of LlamaTouch.
+The above evaluation logic is well encapsulated into the LlamaTouch evaluator.
+To use it for evaluation, an agent only needs to define how to load the task execution traces for each task, as shown in the right side of Figure 2, The evaluator will automatically conduct the evaluation and report metrics such as task completion rate.
+Experiments in §$\S$5 show the faithful evaluation of LlamaTouch.
+4. LlamaTouch Dataset
+4.1. Dataset Construction
+The LlamaTouch dataset consists of a combination of tasks from the previous AITW dataset (Rawles et al., 2023) and self-generated tasks that involve diverse categories and popular applications.
+The inclusion of new tasks in currently popular applications, which have not been covered in previous literature, aims to properly assess the generality and real capabilities of mobile agents.
+Figure 5 demonstrates the workflow for constructing the dataset, where each data sample undergoes the following processes.
+Generate task descriptions.
+Task descriptions are generated by both humans and LLMs using app metadata (e.g., app names, categories, descriptions) from popular apps in the Google Play Store.
+For tasks in the AITW dataset, we sample parts of them after deduplicating similar task descriptions.
+The sampled task descriptions are then validated by humans to avoid duplication, infeasibility, and high complexities that may far exceed the capabilities of human and mobile agents.
+After this process, we sampled 102 tasks from the AITW dataset among 26 unique apps and 394 newly generated tasks among 46 unique apps.
+The generated descriptions cover a variety of application categories such as utilities (e.g., Zoom, Expedia), social media (e.g., Discord, Instagram), and web shopping (e.g., Walmart, Amazon).
+The left-hand side of Figure 5 shows the proportion of these categories.
+We then employ six human annotators, all of whom are authors of this study and experts in smartphone usage, to generate data samples in LlamaTouch through the following two independent annotation stages.
+The image on the lower side of Figure 5 illustrates the outputs of the human annotation process.
+Record UI interaction traces.
+The validated task descriptions are used as guidance for recording UI interaction traces.
+Tasks sampled from the AITW dataset are also required to go through this process as they lack view hierarchies.
+Given a task description, human annotators interact with mobile apps through our developed UI Interaction Recorder.
+This tool is built on top of mobile emulators or realistic smartphones and displays the graphical user interface to users, allowing them to operate it like normal smartphones.
+Specifically, human annotators are asked to complete a task in the simplest manner and to avoid redundant operations, as in (Rawles et al., 2023).
+As shown in Figure 5, the task “Empty the shopping cart on bestbuy” requires five continuous click actions to complete.
+The recorder captures actions, VHs, activities, and screenshots, which collectively form the UI trajectory.
+Annotate essential states.
+The recorded UI interaction traces, along with task descriptions, are then processed by human annotators to identify essential states.
+To help annotators better understand the application state transformation and simplify the annotation process, we developed an essential state annotation system.
+This system displays the entire UI interaction trace, with potentially significant UI components shown with numeric indices in each screenshot.
+Annotators are asked to identify the most significant, identifiable states that represent key milestones during task completion.
+They annotate these states (their numeric indices) with the proper essential state primitives we proposed in §$\S$3.2.
+These essential states, together with recorded UI interaction traces, form the final LlamaTouch dataset.
+During the annotation process, to ensure data reliability, anything that a single human annotator cannot decide on is annotated based on a consensus reached by three or more annotators.
+Overall, this dataset includes 496 tasks, covering 57 unique Android applications with diverse task complexities.
+4.2. Dataset Statistics
+In this section, we present statistics of the LlamaTouch dataset.
+Table 3 quantifies task complexities by showing the average steps (actions) required to complete a task.
+Tasks from AITW (Rawles et al., 2023) are slightly more complex than those generated by LlamaTouch, with average steps of 7.35 versus 5.67, respectively.
+Overall, the average number of steps to complete a task is nearly 7, with task complexities ranging from 2 to 42 steps.
+Table 3. LlamaTouch dataset statistics and task complexities measured by the average steps (actions) to complete a task.
+Category # Task # Apps Avg. Steps
+AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
+Generated 394 46 5.67 (3-42)
+Total 496 57 7.01 (2-42)
+We further analyze action types in the ground-truth dataset.
+Table 4 shows the statistics of actions contributing to the dataset.
+In summary, all 496 tasks involve click actions.
+Out of these, 292 tasks involve swipe actions, and 147 tasks involve input actions.
+A few tasks also use press home and back for navigation.
+By dividing the total number of actions by the number of tasks involving these actions, we observe an average of more than 4 click actions per data sample.
+The average number of all other actions is less than 2.
+Table 4. Action statistics in LlamaTouch dataset.
+Action # Action # Tasks W/ Action Mean/Stddev
+Click 2,192 496 4.42/2.51
+Swipe 376 292 1.29/0.9
+Input Text 173 147 1.18/0.55
+Press Home 44 43 1.02/0.15
+Press Back 6 5 1.2/0.45
+Table 5. Statistics of annotated essential states.
+Type Essential State (ES) # ES # Tasks W/ ES Mean/Stddev
+Exact Match UI Component 698 418 1.67/0.92
+Activity 490 442 1.11/0.32
+Action: Click 98 93 1.05/0.23
+Action: Type 1 1 1/0
+System: Install 7 7 1/0
+System: Uninstall 3 3 1/0
+Fuzzy Match UI Component 51 44 1.16/0.43
+Screen Info 31 31 1/0
+Statistics in Table 5 detail the annotated essential states.
+We find that the essential states of 441 tasks can be presented on only one UI representation, indicating that most tasks check only the final states after task completion.
+There are 53 tasks and 2 tasks with essential states presented on two and three UI representations, respectively.
+Among the annotated essential states, exact UI component match accounts for 51% of all essential states, followed by exact activity match at 36%.
+More than 84% of tasks (418 out of 496) have at least one of these two types of essential states.
+Fuzzy match is also important for the evaluation process: 44 tasks have 51 fuzzy UI component match in total; fuzzy screen info match occur in 31 tasks.
+In summary, all types of essential states construct the dataset, which significantly contributes to the faithful evaluation of mobile UI task automation, as we will show in the following section.
 5. Evaluation
-4.1. Dataset
-Table 3. LlamaTouch’s dataset statistics.
-Task Category # Tasks # Apps Avg. Task Steps
-Tasks from AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
-LlamaTouch Generated 393 46 5.67 (3-42)
-Total 495 57 7.01 (2-42)
-LlamaTouch uses a combination of tasks in the previous AITW dataset (Rawles et al., 2023) and self-constructed ones for evaluation.
-Table 3 shows the statistics of this dataset.
-We sampled 102 tasks from AITW (Rawles et al., 2023), one of the most popular datasets for mobile device control.
-By observing the limited number of evaluated applications, we generated 395 tasks in 46 applications.
-Overall, this dataset includes 495 tasks, covering 57 popular Android applications with different difficulty levels (measured by the number of steps to finish a task).
-Every single data entry consists of three types of data.
-  • A task description in natural language, e.g., empty the shopping cart on bestbuy.
-  • A sequence of raw UI representations when recording these tasks. Each raw UI representation contains a screenshot, VH, activity name, and action performed on the current screen.
-  • The annotated essential states by human annotators.
 5.1. Experiment Setup
-Mobile environment.
+Mobile environments.
 LlamaTouch primarily utilizes an x86-64 Android emulator (Google, 2023) as the mobile environment for task execution.
-The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server with Ubuntu 18.04 OS.
-Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) will be tested on a real Google Pixel 5 device with the Android 14 OS.
+The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server running Ubuntu 18.04 OS.
+Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) are tested on a real Google Pixel 5 device with Android 14 OS.
+Task setup.
+In this study, tasks are set up in two stages.
+First, we construct an Android system image for the Android emulator, which provides a concrete environment where partial task states are prepared.
+This ensures a controlled and replicable starting point for the tasks.
+Second, we develop setup scripts using the Android UIAutomator2 library (uia, 2024) to manage complex setup processes with specific constraints (e.g., require manual login or run on real-world smartphones).
+These scripts automate the initialization of the environment, ensuring tasks are ready for execution with minimal manual intervention.
 Mobile agents. We selected four mobile agents that cover diverse types of brains, including supervised learning models, LLM, and large multi-modality models.
-  • AutoUI (Zhan and Zhang, 2023) that employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and t5 (Raffel et al., 2020) with its variants as the brain for decision making.
-It takes pixel-level screenshots and task descriptions as the input.
-  • AutoDroid (Wen et al., 2023) that uses LLMs for device control.
-It first takes textual screen VH and simplifies it to explicit, readable HTML representation.
-Then, the simplified HTML representation will be processed by LLMs and generate the corresponding action.
+  • Auto-UI (Zhan and Zhang, 2023) employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and T5 (Raffel et al., 2020) variants as the brain for decision-making.
+It takes pixel-level screenshots and task descriptions as input.
+  • AutoDroid (Wen et al., 2024) uses LLMs for device control.
+It first takes textual screen VH and simplifies it to an explicit, readable HTML representation.
+The simplified HTML representation is then processed by LLMs to generate the corresponding action.
 We selected GPT-4 (gpt-4-0125-preview) (Achiam et al., 2023) as the backend of AutoDroid.
-  • AppAgent (Yang et al., 2023) that utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch.
-In our evaluation, we used the version AppAgent without document pre-exploration for simplicity.
-  • CoCo-Agent (Ma et al., 2024) that uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
-All actions generated by these agents are applied to LlamaTouch’s Android environment through one adaption layer.
+  • AppAgent (Yang et al., 2023) utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch.
+In our evaluation, we used the AppAgent version without document pre-exploration for simplicity.
+  • CoCo-Agent (Ma et al., 2024) uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
 Evaluation methodologies. We used the following methodologies for evaluating whether an agent completes a task.
-  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023).
+  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2024; Zhan and Zhang, 2023).
 It compares the agent-generated action sequences in real-world environments with ground-truth action sequences in the dataset.
@@
 The results of human validation are treated as ground truth for comparing the accuracy of the three evaluation approaches.
-Metrics.
-We primarily compare the (1) end-to-end task completion rate (TCR) of different agents and (2) the accuracy of different evaluation approaches, taking human validation as the ground-truth result.
-In step-wise action match and LCS-based action match, a task is completed only when the two action sequences are identical or the ground-truth one serves as the subsequence of the agent-generated action sequence, respectively.
-In LlamaTouch, a task is treated as completed when the agent execution trace passes through all essential states.
-For accuracy, a task is accurately evaluated only if the evaluation result matches the human validation result.
-The accuracy of an evaluation approach denotes the proportion of tasks correctly evaluated out of all tasks.
+5.2. Metrics
+We primarily compare (1) the end-to-end task completion rate (TCR) of different mobile agents and (2) the accuracy of different evaluation approaches, using human validation results as the ground truth.
+End-to-end TCR.
+Executing tasks in realistic mobile environments results in two outcomes: task completion and task non-completion.
+In step-wise action match, a task is considered completed only when the two action sequences are identical.
+LCS-based action match assesses whether a ground-truth action sequence is a subsequence of the agent-generated action sequence; if so, the task is evaluated as completed.
+In LlamaTouch, a task is considered completed when the agent execution trace passes through all essential states in the ground-truth dataset, using the proposed algorithm in §$\S$3.3.
+The end-to-end TCR of a certain evaluation method is calculated as the number of tasks evaluated completed divided by the total number of tasks in the dataset.
+Evaluation accuracy.
+In this study, we use human validation results on agent execution traces as the ground truth for task completion.
+Assume $N$ denotes the ground-truth dataset encompassing all tasks.
+For task $n$ in $N$, we use $H_{n}$ to represent the human validation result based on the agent execution trace, where $H_{n}$ can be either “True” or “False”.
+When evaluating the same agent execution trace using a specific evaluation method, the outcome is represented by $E_{n}$, which can also be “True” or “False”.
+Therefore, the accuracy formula is defined as:
+(1) $\text{accuracy}=\frac{\sum_{n\in N}\left(\delta_{E_{n},H_{n}}\right)}{|N|}$
+The formulation indicates that a task is accurately evaluated only when the evaluation result $E_{n}$ matches the human validation result $H_{n}$.
+Consequently, the accuracy of an evaluation approach is defined as the proportion of correctly evaluated tasks to the total number of tasks in the dataset $|N|$.
 5.3. Task Completion Rate and Accuracy
+Table 6 presents the end-to-end TCR and accuracy of different evaluation designs.
+All three approaches achieve more than 90% accuracy on average.
+However, the step-wise action match and LCS-based action match fall short in recognizing tasks correctly executed by mobile agents, resulting in nearly 0% TCR.
+The high accuracy rates of these two approaches are attributed to the large portion of incomplete tasks; only 6% of tasks are deemed completed upon human validation.
+Compared to action match on static datasets, the TCR evaluated by LlamaTouch is 8.67%, which closely aligns with the results of human validation.
+To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results pervasive in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks.
+Table 7 displays the number of tasks completed by agents and the evaluation accuracy for these tasks.
+Overall, agents successfully completed 30 tasks on average.
+Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy.
+This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets, primarily because a single static action sequence is difficult to match with multiple possible paths to task completion in real-world environments.
+Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods.
+The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
 Table 6. End-to-end task completion rate (TCR %) and accuracy (Acc. %) of different evaluation approaches of all tasks.
@@
 Average 0.00 93.86 0.15 93.81 8.67 94.93 6.14
-Table 4 presents the end-to-end TCR and accuracy of different evaluation designs.
-All three approaches achieve more than 90% accuracy on average.
-However, the step-wise action match and LCS-based action match fall short in recognizing tasks correctly executed by mobile agents, resulting in nearly 0% TCR.
-The high accuracy rates of these two approaches are attributed to the large portion of incomplete tasks; only 6% of tasks are deemed completed upon human validation.
-Compared to action match on static datasets, the TCR evaluated by LlamaTouch is 8.67%, which closely aligns with the results of human validation.
-To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results that pervasively occurred in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks.
-Table 5 displays the number of tasks completed by agents and the evaluation accuracy for these tasks.
-Overall, agents successfully completed 30 tasks on average.
-Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy.
-This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets.
-Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods.
-The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
 Table 7. Accuracy (Acc. %) of different evaluation approaches among all successful tasks in human validation.
@@
 LlamaTouch Human
-Acc. Acc. Acc. # success tasks
+Acc. Acc. Acc. # success
 AutoUI 0.00 0.00 77.78 9
@@
 Average 0.00 0.76 78.91 30
-$4.$ Ablation Study
+5.4. Ablation Study
+LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributed to the annotation and implementation of various essential state primitives.
+In this section, we evaluate the effectiveness of two types of match designs in LlamaTouch: fuzzy match and exact match.
+We aggregate the execution traces of all mobile agents evaluated in §$\S$5.3.
+We evaluate the system by first disabling exact match and fuzzy match separately, and then enabling different primitives one at a time.
+We present the results for (1) all tasks, (2) tasks in AITW (Rawles et al., 2023), and (3) new tasks generated in LlamaTouch.
+The results for task completion rate and accuracy are shown in Table 8.
+Exact match ensures key information on two screens is identical.
+As shown in the results, exact match significantly contributes to improving evaluation accuracy.
+For example, without all exact match primitives, LlamaTouch’s evaluation accuracy on all tasks drops from 95% to 19%.
+Among all exact match primitives, activity match greatly improves evaluation accuracy.
+This is due to its simple yet efficient ability to locate functional application screens and the large proportion of annotated activity primitives, comprising 35% of all primitives in our dataset.
+UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen.
+Exact match for action and system state only improves accuracy slightly because only a few tasks are annotated with these primitives.
+Fuzzy match does not show a notable improvement in evaluation accuracy, despite the presence of 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset.
+Few tasks with these primitives can be completed by the four agents we evaluated, leading to only a slight improvement.
+Although primitives for fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world.
+With more performant mobile agents in the future, they will be further explored and evaluated.
 Table 8. Ablation study on different essential state primitives.
-Evaluation design All tasks AITW tasks LlamaTouch tasks
+Evaluation design All tasks AITW Generated
 TCR Acc. TCR Acc. TCR Acc.
@@
 + textbox fuzzy match 8.97 94.83 18.69 88.69 6.45 96.43
-LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributing to the annotation and implementation of different essential state primitives.
-In this section, we evaluate the effectiveness of two types of match design in LlamaTouch: fuzzy match and exact match.
-We aggregate the execution traces of all mobile agents evaluated in $\S$§4.3.
-As for evaluation implementation, we disable exact match and fuzzy match separately, and then enable different primitives once per time.
-We present the results on (1) all tasks; (2) tasks in AITW (Rawles et al., 2023); and (3) tasks generated in LlamaTouch.
-The results of task completion rate and accuracy are shown in Table 6.
-Exact match ensures key information in two screens are identical.
-As shown in the results, exact match contributes a significant portion of evaluation accuracy improvement.
-For example, without all exact match primitives, the evaluation accuracy of LlamaTouch on all tasks drops from 95% to 19%.
-Among all exact match primitives, activity match highly improves the validation accuracy.
-This is not only attributed to its simple yet efficient ability to locate functional application screens, but also the large portion of annotated activity primitives, i.e., 35% out of all primitives in our dataset.
-UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen.
-Action and system state exact match only improve a little accuracy as only a few tasks are annotated with these primitives.
-Fuzzy match does not show notable improvement in evaluation accuracy, although there are 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset.
-Seldom of tasks with these primitives can be completed by the four agents we evaluated, leading to a slight improvement.
-Although primitives about fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world.
-With more performant mobile agents in the future, they will be further explored and evaluated.
 5.5. Absolute Capabilities of Mobile Agents
 In this section, we present the absolute capabilities of different mobile agents in mobile UI automation tasks.
-We first categorize tasks according to their sources: AITW and our self-constructed dataset.
-This aims to reveal whether an agent that has previously learned on a well-established dataset has the capability to adopt new applications/tasks.
-Second, we categorize tasks according to their difficulty, measured by the number of steps required to complete a task in the ground-truth dataset.
-Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 <<< steps $\leq$ 8), and high (step >>> 8).
-Table 7. The comparison of task completion rate categorized by task sources (e.g., from the previous dataset AITW or from LlamaTouch’s generated dataset).
+First, we categorize tasks according to their sources: AITW and our self-generated dataset.
+This aims to determine whether an agent that has previously learned on a well-established dataset can adapt to new applications/tasks.
+Second, we categorize tasks according to their complexities, measured by the number of steps required to complete a task in the ground-truth dataset.
+Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 $<$ steps $\leq$ 8), and high (step $>$ 8).
+Performance across datasets.
+Among the four evaluated agents, AutoUI and CoCo-Agent were previously trained on AITW.
+AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively.
+We separately show the end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to demonstrate their performance generalization to new scenarios (i.e., new tasks and new applications).
+Results in Table 9 indicate that AutoDroid and AppAgent also achieve higher TCR in AITW tasks compared to LlamaTouch’s generated tasks.
+This is mainly attributed to the task complexity gap between the two datasets.
+AutoUI and CoCo-Agent achieve a 12.75% TCR on tasks in AITW.
+However, for generated tasks in LlamaTouch, both perform poorly, with only 2.29% and 2.31% TCR, respectively.
+The TCR gap between datasets reveals their lack of capability to adapt to previously unseen tasks.
+Considering the vast number of real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
+Table 9. The comparison of task completion rate categorized by task sources (i.e., from AITW and LlamaTouch generated tasks).
 Agent End-to-end TCR
 Overall AITW Generated
-AutoUI 4.44 12.75 2.29
+Auto-UI 4.44 12.75 2.29
 AutoDroid 14.84 22.77 12.79
@@
 CoCo-Agent 4.47 12.75 2.31
-Performance among datasets.
-Among the four evaluated agents, AutoUI and CoCo-Agent are previously trained on AITW.
-AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively.
-We separately show end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to show their performance generalization to new scenarios (i.e., new tasks and new applications).
-Results in Table 7 show that AutoDroid and AppAgent also show higher TCR in AITW tasks than LlamaTouch’s generated tasks.
-This mainly attributes to the task complexity gap between the two datasets.
-AutoUI and CoCo-Agent achieve 12.75% TCR on tasks in AITW.
-However, for generated tasks in LlamaTouch, both of them operate poorly: with only 2.29% and 2.31% TCR, respectively.
-The TCR gap between datasets reveals their lack of capability to adapt to unseen tasks before.
-Considering massive real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
 Performance under different task complexity.
-We categorize tasks into three difficulty levels according to the step number required to complete the task in the datasets.
+We categorize tasks into three difficulty levels based on the number of steps required to complete them in the datasets.
 The results are shown in Table 10.
@@
 There is a significant drop in TCR when task complexity increases.
-AutoDroid outperforms the other three agents for tasks under all difficulties.
-AppAgent achieves a similar TCR among all tasks as AutoDroid.
-We deem that the knowledgable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations.
+AutoDroid outperforms the other three agents for tasks at all difficulty levels.
+AppAgent achieves a TCR similar to AutoDroid across all tasks.
+We believe that the knowledgeable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations.
 Therefore, agents based on GPT-4/4V achieve high TCR.
-However, there is still huge space for mobile agents to improve their capabilities in UI automation tasks.
+However, there is still considerable room for mobile agents to improve their capabilities in mobile UI task automation.
 Table 10.
 The comparison of task completion rate between mobile agents.
-Tasks are categorized to different difficulty levels according to the number of steps required to finished it in the ground-truth dataset.
+Tasks are categorized into different difficulty levels according to the number of steps required to finish it in the ground-truth dataset.
 Agent End-to-end TCR
 […]
-AutoUI 4.44 4.95 4.19 4.82
+Auto-UI 4.44 4.95 4.19 4.82
 AutoDroid 14.84 27.00 12.94 7.32
@@
 Supporting WebView-based apps.
-One major limitation of LlamaTouch is that its evaluation process requires Android’s view hierarchy to empower approximate UI layout matching and accurate UI component matching ($\S$§3.2).
-While some Android applications built with WebView (Google, 2024a) are unable to access their view hierarchy, making LlamaTouch cannot evaluate agent performance on these applications.
-However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing view hierarchy.
+One limitation of LlamaTouch is that its evaluation process requires Android’s VH to empower approximate UI layout matching and accurate UI component matching (§$\S$3.2).
+Some Android applications built with WebView (Google, 2024a) are unable to access their VH, preventing LlamaTouch from evaluating agent performance on these applications.
+However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing VH.
 When evaluating WebView-based apps are required, more advantaged techniques, such as screen similarity detection (Wu et al., 2023b, a), OCR/model-powered screen element recognition (Bunian et al., 2021; Xie et al., 2020), should be incorporated to retrofit LlamaTouch.
@@
 Biases in identifying essential states.
-LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on pre-defined task execution traces.
+LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on predefined UI interaction traces.
 This process, however, could involve biases and leads to inaccuracy due to potentially limited knowledge in application execution and unseen circumstances.
-For example, given a task “Delete YouTube in the Google Play Store”, the pre-defined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution.
+For example, given a task “Delete YouTube in the Google Play Store”, the predefined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution.
 However, they may ignore that this application might be not installed at all.
@@
 7. Conclusion
-In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with faithfulness and scalability in UI automation tasks.
-LlamaTouch enables mobile agents to be tested on real mobile devices.
-At the evaluation stage, it only matches the task execution traces with annotated essential states.
-LlamaTouch well tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches.
-It achieves high evaluation accuracy like human validation, while preserving the scalability of evaluating on static datasets.
-By putting task execution on real mobile devices, we also reveal the limited capabilities of current mobile agents in UI automation tasks.
+In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with both faithfulness and scalability in mobile UI task automation.
+LlamaTouch enables mobile agents to be tested on realistic mobile environments.
+At the evaluation stage, it matches the task execution traces with annotated essential states.
+LlamaTouch tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches.
+It achieves high evaluation accuracy, which is comparable to human validation, while preserving the scalability of evaluating on static datasets.
+By conducting task execution to real-world mobile devices, we also reveal the limited capabilities of current mobile agents in mobile UI automation tasks.
+Acknowledgements.
+This work was supported by National Key R&D Program of China (No.2021ZD0113001), NSFC (No.62102045 and No.62272261), and China Institute of IoT (Wuxi).
 Appendix A Appendix
-A.1. APIs Provided by LlamaTouch
+A.1. APIs Provided by AgentEnv
 LlamaTouch provides three categories of APIs for mobile agent integration and UI automation task execution. APIs, parameters, and their return values are listed in Table 11.